OpenStudy (anonymous):

Steve buys last years best selling novel, in hardcover, for $13.00. This is with a 35% discount from the original price. What was the original price of the novel?

OpenStudy (anonymous):

13 = 65% of original price 13/65 = 1% = 1/5 1/5 x 100 =$20

OpenStudy (anonymous):

$20

OpenStudy (anonymous):

gianfraco can you please explain how you got $20? im really confused on this one

OpenStudy (anonymous):

Yes of course

OpenStudy (anonymous):

If something is bought for at a 35% discount it is equal to (100% - 35%) of the original cost. So it is worth 65% of the original.

OpenStudy (anonymous):

ah thank you! (i really hate summer break...)

OpenStudy (anonymous):

So now by dividing $13 by 65 you get to see the value of 1% of the originial price. 13/65 = 1/5 or 0.2 or even better 20c.

OpenStudy (anonymous):

Now multiply 20c by 100 to get the original price of $20

OpenStudy (anonymous):

x = 100 13 = 65 x = 13*100/65 = $20

OpenStudy (anonymous):

Enjoy your summer life is too short to worry about mathematics on a sunny day!

OpenStudy (anonymous):

Well thanks but I really love algebra and geometry set me back so I'm working hard this smmer so I'll be ready for alegbra II

OpenStudy (anonymous):

Good luck with it then. I hope I can help you again!